# PrivATE: Differentially Private Confidence Intervals for Average Treatment Effects

### v1

@@ B.1 Proofs of the supporting lemmas
Lemma 4.2. The IF of the AIPW learner is given by the influence function of the final mean estimation in the second step. Specifically, the influence stemming from training the nuisance estimators $\pi$ and $\mu$ can be neglected during the privatization step if the estimators are sufficiently stable.
@@
Our proof makes use of results for two-stage influence functions [69] and influence functions of two-stage Neyman-orthogonal learners [60].
@@
where $M=-\int\frac{\partial}{\partial\xi}\psi_{2}(\tilde{z}^{(2)},h(\tilde{z}^{(1)},S(P)),\xi)\mathop{}\!\mathrm{d}F(\tilde{z})$ and $\mathrm{IF}(z,S,P)$ where denotes the IF of the first-stage estimation [69].
@@
To show (i), we observe that $\psi_{2}=\frac{1}{n}\sum_{i=1}^{n}\Gamma_{\hat{\eta}}(z_{i})-\tau$. Therefore, we directly get the desired equality $M=1$. To show (ii), we observe that, for the AIPW estimator, the integral equals $\mathbb{E}[\frac{\partial}{\partial\nu}h(\tilde{z}^{(1)},\nu)]=\mathbb{E}[\frac{\partial}{\partial\eta}\Gamma_{\hat{\eta}}(z)]$, where $\eta=(\hat{\mu},\hat{\pi})$. Similar to Schröder et al. [60], we now observe that, due to the Neyman-orthogonality of the AIPW estimator [e.g., 64] and the stability of the nuisance estimators, we have $\mathbb{E}[\frac{\partial}{\partial\eta}\Gamma_{\hat{\eta}}(z)]=0$, which yields the desired result. ∎
@@ B.2 Proofs of the main theorems
Theorem 4.3. Let $z:=(a,x,y)$ define a data sample following the joint distribution $\mathcal{Z}$ and $\hat{\eta}=(\hat{\mu},\hat{\pi})$ the estimated nuisance functions. Furthermore, let $D$ be the training dataset with $|D|$ = $n$. Define
@@
[…]
@@
[…]
@@
[…]
[see 50]. However, similar to the global sensitivity, the smooth sensitivity is extremely expensive or even infeasible to calculate, depending on the machine learning model employed for nuisance estimation. Therefore, we follow Avella-Medina [5] and aim to upper bound the smooth sensitivity by the appropriately scaled gross-error sensitivity of the AIPW estimator.
@@
B.2.2 Proof of Theorem 4.5
Theorem 4.5. Let $z:=(a,x,y)$ define a data sample following the joint distribution $\mathcal{Z}$, and let $\hat{\eta}=(\hat{\mu},\hat{\pi})$ be the estimated nuisance functions. Furthermore, let $D$ be the training dataset with $|D|$ = $n$. For $U\sim\mathcal{N}(0,1)$, we define
@@
[…]
@@
Theorem 4.6. Let $z:=(a,x,y)$ define a data sample following the joint distribution $\mathcal{Z}$, and let $\hat{\eta}=(\hat{\mu},\hat{\pi})$ be the estimated nuisance functions. We define the privacy budget $(\varepsilon,\delta)$ and the sub-budgets $\varepsilon_{1},\varepsilon_{2},\delta_{1},\delta_{2}$, for ATE and variance estimation under DP such that $\varepsilon_{1}+\varepsilon_{2}=\varepsilon$ and $\delta_{1}+\delta_{2}=\delta$. Furthermore, let $D$ be the training dataset with $|D|$ = $n$, and let $\hat{\tau}_{\textrm{{DP}}}$ denote the $(\varepsilon_{1},\delta_{1})$-differentially private ATE estimate and $\hat{\sigma}_{\textrm{{DP}}}$ denote the $(\varepsilon_{2},\delta_{2})$-differentially private variance estimate. Then, the valid and $(\varepsilon,\delta)$-differentially private CI is given by
@@ C.1.1 Synthetic data
We generate our datasets similar to Oprescu et al. [53]:
@@ C.2 Implementation details
The PrivATE framework includes the calculation of a supremum over a bounded data space. We implemented this maximization problem using L-BFGS-B [14], a limited-memory Broyden–Fletcher–Goldfarb–Shanno algorithm for solving nonlinear optimization problems with bounded variables. We use default hyperparameters and use $n=10$ random starting points for every optimization.
@@ Appendix E Extension to interventional RCT data
We now demonstrate the applicability of PrivATE to experimental data. In doing so, we can now additionally compare our method against the private differences of means method by Guha and Reiter [29], yet which is restricted to experimental data (and thus not applicable to the observational data used throughout our main paper). Again, we evaluate PrivATE using simulated and real-world RCT data.
@@ E.2 Results
Figure 6: Visualization of the CIs on RCT data. The CIs are reported for both PrivATE and the difference in means method by Guha and Reiter [29]. For both, we use the NN base learner across different runs with privacy budgets $\varepsilon=0.5$ and $\delta=10^{-5}$. $\Rightarrow$ The intervals of the difference in means method are significantly wider than the ones of PrivATE. Our method returns more informative intervals and is thus preferred.

### v2

@@ B.1 Proofs of the supporting lemmas
Lemma 4.2. We make the following assumptions on the nuisance functions: The nuisance functions are (i) bounded, (ii) estimated at rates $n^{-\beta_{\mu}}$ and $n^{-\beta_{\pi}}$ with $\beta_{\mu}+\beta_{\pi}\geq\frac{1}{2}$, and (iii) in a local neighborhood of the true nuisance functions, i.e., there exists $\lambda_{n}$ decreasing in the sample size $n$, s.t. $\rVert\hat{\eta}-\eta_{0}\lVert{\infty}\leq\lambda_{n}$. Then, the IF of the AIPW learner is dominated by the influence function of the final mean estimation in the second step. The influence stemming from training the nuisance estimators $\hat{\pi}$ and $\hat{\mu}$ is negligible in the privatization step.
@@
Our proof makes use of results for two-stage influence functions (Zhelonkin et al., 2012).
@@
[…]
@@
[…]
$\Gamma_{\hat{\eta}}(z)=\Gamma_{\eta_{0}}(z)+\frac{\partial}{\partial\eta}\Gamma_{\eta_{0}}(z)(\hat{\eta}-\eta_{0})+R_{2},$ (28)
where $R_{2}$ includes higher order terms, and thus
$\mathbb{E}[\frac{\partial}{\partial\eta}\Gamma_{\hat{\eta}}(z)]=\mathbb{E}[\frac{\partial}{\partial\eta}R_{2}].$ (29)
Recall that by the local neighborhood assumption there exists $\lambda_{n}$ small and decreasing in $n$, such that $\rVert\hat{\eta}-\eta_{0}\lVert{\infty}\leq\lambda_{n}$, and thus as well $R_{2}<<\lambda_{n}$. Note that $R_{2}$ depends on the Hessian matrix of $\Gamma_{\eta}$ with respect to the estimated nuisances. If the propensity is known, the Hessian vanishes by straightforward computation, and thus $\mathbb{E}[\frac{\partial}{\partial\eta}\Gamma_{\hat{\eta}}(z)]=0$. If the propensity is unknown, we make use of the local neighborhood property in (Chernozhukov et al., 2018), resulting in $\mathbb{E}[\frac{\partial}{\partial\eta}\Gamma_{\hat{\eta}}(z)]=\mathbb{E}[\frac{\partial}{\partial\eta}R_{2}]=0$. Overall, the IF of the AIPW estimator is thus dominated by the influence function of the final mean estimation in the second step. ∎
@@ B.2 Proofs of the main theorems
[…]
@@
Our proof builds upon the idea of upper-bounding the global sensitivity $\Delta_{2}(f)$ required for privatization with that Gaussian mechanism (Definition 3.2) by the so-called smooth-sensitivity (Nissim et al., 2007). Specifically, for the $\xi$-smooth sensitivity $\mathit{SS}_{\xi}(f,D)$ of the estimator $f$ with $\xi=\frac{\varepsilon}{4(d+2\log(2/\delta))}$, it holds that
$f^{\textrm{{DP}}}_{D}=f_{D}+\frac{5\sqrt{2\log(2/\delta)}}{\varepsilon}\mathit{SS}_{\xi}(f,D)\cdot U,$ (33)
@@
$\mathit{SS}_{\xi}(f,D):=\sup_{D^{{}^{\prime}}}\{\exp(-\xi d_{\mathrm{H}}(D,D^{{}^{\prime}}))LS(f,D^{{}^{\prime}})\mid D^{{}^{\prime}}\in\mathcal{Z}^{n}\},$ (34)
@@
$\mathit{LS}(f,D):=\sup_{D^{{}^{\prime}}}\left\{\lVert f_{D}-f_{D^{{}^{\prime}}}\rVert\;\Big|\;d_{\mathrm{H}}(D,D^{{}^{\prime}})=1\right\}$ (35)
(see Nissim et al., 2007). However, similar to the global sensitivity, the smooth sensitivity is extremely expensive or even infeasible to calculate, depending on the machine learning model employed for nuisance estimation. Therefore, we follow Avella-Medina (2021) and upper bound the smooth sensitivity by the appropriately scaled gross-error sensitivity of the AIPW estimator.
@@
The influence function is differentiable with respect to $\hat{\tau}$ with derivative $\frac{\partial}{\partial\hat{\tau}}\mathrm{IF}^{\textrm{{AIPW}}}(z,\tau_{0};P)=1$. Furthermore, since we assume the data to stem from a bounded domain, there exists a constant $K\in\mathbb{R}$ such that, for the gross-error sensitivity of $\hat{\tau}$, one has $\gamma_{\tau}\leq K$. As a result, we can upper bound the smooth sensitivity of the AIPW estimator by $\mathit{SS}_{\xi}(\hat{\tau},D)\leq\frac{\sqrt{\log{(n)}}}{n}\gamma_{\tau}$ (Avella-Medina, 2021), leading to the desired result in Theorem 4.3. ∎
B.2.2 Proof of Theorem 4.5
@@
We proceed in the same manner as in the proof of Theorem 4.3. We observe that $\hat{\sigma}_{\textrm{{AIPW}}}^{2}$ solves $\sum_{i=1}^{n}\mathrm{IF}^{\sigma}(z_{i},\sigma_{0};P)=0$ with $\frac{\partial}{\partial\hat{\sigma}^{2}}\mathrm{IF}^{\sigma}(z,\sigma_{0};P)=1$. Again, since we assume the data to stem from a bounded domain, there exists a constant $K\in\mathbb{R}$ such that, for the gross-error sensitivity of $\hat{\sigma}^{2}$, we have $\gamma_{\sigma}:=\sup_{z\in\mathcal{Z}}\big\lVert\mathrm{IF}^{\sigma}(z,\sigma_{0},P)\big\rVert\leq K$. Then, with Theorem 1 in Avella-Medina (2021), we can upper bound the smooth sensitivity $\mathit{SS}_{\xi}(\hat{\sigma}^{2},D)$ by $\frac{\sqrt{\log{(n)}}}{n}\gamma_{\sigma}$. As a result, we get that the quantity
@@
[…]
@@ C.1.1 Synthetic data
We generate our datasets similar to Oprescu et al. (2019):
@@ C.2 Implementation details
The PrivATE framework includes the calculation of a supremum over a bounded data space. We implemented this maximization problem using L-BFGS-B (Byrd et al., 1995), a limited-memory Broyden–Fletcher–Goldfarb–Shanno algorithm for solving nonlinear optimization problems with bounded variables. We use default hyperparameters and use $n=10$ random starting points for every optimization.
@@ Appendix E Extension to interventional RCT data
We now demonstrate the applicability of PrivATE to experimental data. In doing so, we can now additionally compare our method against the private differences of means method by Guha and Reiter (2024), yet which is restricted to experimental data (and thus not applicable to the observational data used throughout our main paper). Again, we evaluate PrivATE using simulated and real-world RCT data.
@@ E.2 Results
Figure 6: Visualization of the CIs on RCT data. The CIs are reported for both PrivATE and the difference in means method by Guha and Reiter (2024). For both, we use the $N$ base learner across different runs with privacy budgets $\varepsilon=0.5$ and $\delta=10^{-5}$. $\Rightarrow$ The intervals of the difference in means method are significantly wider than the ones of PrivATE. Our method returns more informative intervals and is thus preferred.